A grocery store buys 20 pounds of beef for $18.50 and marks it up 35%. What is the price for the customer?
Answer by ewatrrr(24785)


Hi,
20 pounds of beef purchursed for $18.50 wholesale marked up 35%
18.50 + .35*18.50 = 18.50 +6.48 = 24.98/20lbs
OR 1.25/lb would be the cost to the customer